the regression equation of x and y is 3y-5x+108=0.if the mean value of y is 44 and the variance of x were 9/16 th of the variance of y .find the mean value of x ang correlation coefficient
1
Expert's answer
2020-05-13T18:25:57-0400

"X\text{ is independent variable}, Y\text{ is a dependent variable}.\\
y=mx+b\\
m=\frac{\overline{xy}-\overline{x}\ \overline{y}}{\overline{x}^2-(\overline{x})^2}\\
b=\overline{y}-m\overline{x}\\
3y-5x+108=0\\
y=\frac{5}{3}x-36\\
m=\frac{5}{3}\\
b=-36\\
-36=44-\frac{5}{3}\overline{x}\\
\overline{x}=48\\
m=\rho\frac{S_{Y}}{S_{X}}\text{ where }\rho \text{ is a correlation coefficient and }\\
S_{Y}, S_{X} \text{ are sample standard deviations of } X \text{ and } Y\\
\text{respectively}.\\
S_{X}=\frac{9}{16}S_{Y}\\
\frac{S_{Y}}{S_{X}}=\frac{16}{9}\\
\frac{5}{3}=\rho\frac{16}{9}\\
\rho=\frac{\frac{5}{3}}{\frac{16}{9}}=0.9375"